OpenStudy (anonymous):

If sin θ = 21/29 , use the Pythagorean Identity to find cos θ.

OpenStudy (anonymous):

\[\cos \theta = \sqrt{1 - \sin \theta ^{2}}\]

OpenStudy (anonymous):

i thought i needed to use sin^2+cos^2=1?

OpenStudy (anonymous):

oh wait thats the same thing isnt it....

OpenStudy (anonymous):

they are the same

OpenStudy (campbell_st):

just draw a triangle  find x using pythagoras then you can find the cos ratio

OpenStudy (anonymous):

i got now thx